OpenStudy (anonymous):

10-xy --- 2 a.6 b.11 c.1 d.2

OpenStudy (anonymous):

What's the whole question?

OpenStudy (anonymous):

its 10 - xy thats over 2

OpenStudy (anonymous):

Evaluate the expression when x = 6 and y = 3.

OpenStudy (anonymous):

yes 10- xy over2

OpenStudy (anonymous):

gotcha so its 10 - 6(3)/2

OpenStudy (anonymous):

\[10-\frac{ xy }{ 2 }\] \[10-\frac{ 6(3) }{ 2 }\] \[10-\frac{ 18 }{ 2 }\] Can you figure out the rest? Just divide 18 / 2 then add 10 to it.

OpenStudy (anonymous):

Do the 6 times 3 first giving you 10-18/2.....this^

OpenStudy (anonymous):

19?

OpenStudy (anonymous):

Sorry, I meant subtract 10 from it.

OpenStudy (anonymous):

Not add.

OpenStudy (anonymous):

9?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

then the 10-9

OpenStudy (anonymous):

18 / 2 = 9. Then you have to subtract 10 from it.

OpenStudy (anonymous):

1

OpenStudy (anonymous):

there you go

OpenStudy (anonymous):

What's 10 - 9? That's your answer.

OpenStudy (anonymous):

10 - 9 = 1. So that's your answer.